# Pareto-optimality of pulses for robust population transfer in a ladder-type qutrit

### v2

@@ II.2 Pulse models
where we subtract by $k_{2}$ in Eq. 7 to ensure that we satisfy the condition $\Delta(0)-\delta=0$, which is naturally satisfied for the other detuning functions we consider. For these detuning functions, the parameters $k_{1}$ and $k_{2}$ are searched for in the aforementioned ranges $k_{1}\in[-0.3,0.3]$ GHz and $k_{2}\in[-0.1,0.1]$ GHz; whereas, the parameters $\gamma_{1}$ and $\gamma_{2}$ are each searched for in the range $(0,8]$. The parameter $k_{1}$ acts as a scaling factor which symmetrically affects at which value the detuning saturates, i.e., the detuning asymptotically approaches ±$\pm k_{1}$ since $\tanh(x)\rightarrow\pm 1$ when |$\absolutevalue{x}$ | is large. However, $\gamma_{1}$ controls how quickly the detuning transitions from negative to positive values, i.e., the steepness of the slope. Small values of $\gamma_{1}$ yield a near-linear detuning and large values yield a sharper transition near $t=0$, with extremely large values producing a step-like detuning. Since Eq. 7 is just Eq. 6 plus an additional term, the effect of these parameters is the same in both detuning functions. Moreover, the additional term in Eq. 7 introduces a dip in the detuning function around $t=0$ and increasing $k_{2}$ makes the dip more pronounced. The parameter $\gamma_{2}$ affects the localization of the dip, with a larger $\gamma_{2}$ yielding a dip more concentrated around $t=0$ and a smaller $\gamma_{2}$ making it spread out over a longer time.
As we have previously introduced the shorthand notations SG and SECH for our two envelope types, we further systematically distinguish our models by referring to Eqs. 4, 5, 6, and 7 as detunings of type 1, type 2, type 3, and type 4, respectively. We then create our model names by agglutinating the SG or SECH suffix to a particular detuning model as shown in Table 1. Thus, in this work, we consider a total of eight models: four with super-Gaussian envelopes and four with hyperbolic secant envelopes. These models each satisfy the conditions $\Omega(0)=\Omega_{0}$, $\Omega(\pm T/2)=0$, $\Delta(0)=\delta$, and are evaluated at pulse durations of $T=50$ ns and $T=200$ ns. These two pulse durations span a range of experimentally relevant gate times in superconducting qubit hardware. Shorter durations are motivated by the need for fast gate operations in near-term devices, whereas longer durations allow greater detuning robustness and reduced leakage into higher excited states. As an example, in transmon-based systems, we have previously shown that sufficiently long pulses driven close to resonance with modest modulation depths can exhibit detuning robustness with a negligible final second excited state population, thereby isolating the qubit subspace despite the transmon’s weak anharmonicity [16].
@@
Model Envelope $[\Omega(t)/\Omega_{0}]$ Detuning $[\Delta(t)-\delta]$
@@
SG2 $\left(\frac{e^{\beta|2t/T|^{n}}-e^{\beta}}{1-e^{\beta}}\right)$ $\sum\limits_{n=1}^{5}k_{n}(2t/T)^{n}$
@@
Table 1: The two-state-inspired models used in this work, each consisting of an amplitude (pulse envelope) function $\Omega(t)$ and detuning function $\Delta(t)$. Each envelope function has a total of three parameters, i.e., $[\Omega_{0},n,\beta\;(\epsilon)]$ for SG (SECH) envelopes. Here, the detuning prefactors are denoted as $k_{1}$, $\cdots$, $k_{5}$, with detuning models of type 1 and 3 only using $k_{1}$, type 4 using $k_{1}$ and $k_{2}$, and type 2 using all 5. The parameters $\gamma_{1}$, and $\gamma_{2}$ only appear in the detuning functions of types 3 and 4. Each model satisfies the conditions $\Omega(0)=\Omega_{0}$, $\Omega(\pm T/2)=0$, and $\Delta(0)=\delta$.
II.3 Pareto-optimization
@@
The conundrum of determining the optimal pulse parameters for our models arises from the fact that a pulse that maximizes detuning robustness $\Delta_{\rm{rob}}$ at a given threshold of $p_{\rm{e}}(T/2)$ will generally not minimize the maximum transient second-excited state population, $\max_{t}(p_{\rm{f}})$, and vice versa. This creates a multiobjective optimization (MOO) problem, where we must balance these competing objectives. To address this trade-off, we apply evolutionary computation methods to identify the best solutions. Techniques such as genetic algorithms and estimation of distribution algorithms have been successfully used in complex optimization problems across various fields, including evacuation planning [47] and robotic control optimization [48]. Genetic algorithms operate by mimicking the principles of natural selection to iteratively evolve a population of candidate solutions toward optimality. The process begins with the generation of an initial population that satisfies predefined bounds and constraints. Each individual in the population is evaluated using an objective function to assign a fitness score. The algorithm then progresses through iterative stages, where parents for the next generation are selected based on their fitness [49].
@@
In solving this MOO, we rely on the Pareto front, which represents the optimal trade-off between objectives in the objective space. Pareto fronts have been used in multiobservable quantum control, where it was shown that the smoothness and convexity of the Pareto front depend on the commutativity of the observables [26]. To find the Pareto front numerically, we use MATLAB’s gamultiobj() function, which implements a controlled, elitist genetic algorithm based on the non-dominated sorting genetic algorithm (NSGA-II) [50].
@@
Building on these general principles, NSGA-II enhances the process with additional layers of sophistication, such as non-dominated sorting and crowding distance measures. After scoring the offspring based on their objective function values and feasibility, the current population and offspring are merged into a single set. This set is sorted into non-dominated fronts and individuals are ranked based on both their Pareto rank and their crowding distance—a measure of solution diversity within a front. Finally, the population is trimmed to its original size, retaining individuals from higher-ranking fronts and preserving diversity. This iterative process continues until convergence criteria are met, yielding an accurate approximation of the Pareto front and enabling efficient exploration of multiobjective optimization landscapes.
III Results
@@
In this section, we present the Pareto fronts for each of our models which optimally minimizes the maximum transient second-state population $\max_{t}(p_{\rm{f}})$ in the range $t\in[-T/2,T/2]$ given that we must simultaneously maximize the detuning robustness at the threshold $p_{\rm{e}}(T/2)\geq 0.99$. These fronts are evaluated for pulse durations of $T=50$ ns and $T=200$ ns. Since this threshold may be reached at multiple points across the detuning (or amplitude) axis, a consistent definition of robustness is essential. We therefore define detuning robustness (or amplitude robustness) as the width of the $p_{\rm{e}}$ curve between the first two points where $p_{\rm{e}}(T/2)=0.99$, provided that $p_{\rm{e}}(T/2)\geq 0.999$ is reached at least once within this interval. We include amplitude robustness in parenthesis, because, even though it was not one of our objectives in our MOO problem, we compute it using the optimal pulse parameters found by the genetic algorithm. However, we do not discuss it until the next section.
@@
Moreover, the Pareto fronts for each of our models at the two pulse durations $T$ are presented in Fig. 2. Here, we observe that the models form two distinct clusters: one for $T=50$ ns pulses (solid lines) and one for $T=200$ ns pulses, with the latter affording us more detuning robustness. For detuning robustness values up to approximately $0.1$ GHz for 50 ns pulses and $0.15$ GHz for 200 ns pulses, all models yield similar $\max_{t}(p_{\rm{f}})$, with the clusters corresponding to respective pulse durations. This suggests that, within these moderate robustness regimes, performance is largely insensitive to the specific choice of pulse model.
@@
We further note that linear detuning functions perform less optimally compared to the other detuning types at higher detuning robustness levels, as evidenced the divergence towards higher $\max_{t}(p_{\rm{f}})$ values at approximately 0.13 GHz for $50$ ns pulses and at approximately 0.22 GHz for $200$ ns pulses. Otherwise, it is generally less obvious how the other models fare relative to each other. Thus, to better contextualize these results, we assign approximate theoretical bounds to these fronts to ascertain how close we are to the limits imposed by adiabaticity.
@@
Figure 2: The maximum transient second excited state population $\max_{t}(p_{\rm{f}})$ vs detuning robustness for various models with pulse durations of 50 ns (solid curves) and 200 ns (dashed curves). The shaded regions (cyan and magenta for 50 ns and 200 ns pulses, respectively) represent the range of data (convex hull) constituting the fronts found from using the Landau-Zener (LZ) formula to approximate the theoretical limit of the detuning robustness for each model. These LZ-based bounds are computed at the threshold condition $p_{\rm{e}}^{\rm{thresh}}=0.99$, but without requiring that $p_{\rm{e}}(T/2)\geq 0.999$ anywhere within the corresponding detuning interval (i.e., between the roots $\delta_{-}$ and $\delta_{+}$, where $p_{\rm{e}}(\delta_{\pm})=0.99$). The orange dotted boundary includes this additional constraint. For the models with nonlinear detuning, $\Delta(t)$ were expanded about $t=0$ and the chirp prefactor (Landau-Zener velocity) of the linear term was used in the LZ formula.
@@
To this end, we use the Landau-Zener (LZ) formula in calculating the detuning robustness at the threshold $p_{\rm{e}}(T/2)\geq 0.99$. The LZ formula provides the probability of a diabatic transition, i.e.,
@@
where $\Gamma=\Omega(t_{0})^{2}/4\hbar|v|$, which quantifies how adiabatic the transition is. Here $t_{0}$ is defined as the time at $\Delta(t)=0$, i.e., $t_{0}=-\delta/v$. In the limit $\Gamma\gg 1$, the transition is adiabatic and if $\Gamma\ll 1$, the transition is nonadiabatic.
The chirp prefactor $v$ of the linear term of the detuning function is referred to as the Landau-Zener velocity and represents the slope of $\Delta(t)$ at the avoided crossing, i.e.,
$v=\left.\frac{d\Delta(t)}{dt}\right|_{t=0}.$ (9)
Importantly, $v$ dominates regardless of the detuning function. Specifically, for detunings of type 1 and 2, $v=2k_{1}/T$, whereas for detunings of type 3 and 4, $v=2k_{1}\gamma_{1}/T$ In terms of $v$ and considering the threshold value of the first excited state population we use for calculating the detuning robustness, i.e., $p_{\rm{e}}^{\rm{thresh}}=0.99$, we impose the condition
@@
Here, we subtract from unity as Eq. 8 describes the probability of remaining in $|\rm{g}\rangle$ at the level crossing, and we further assume that $|\rm{f}\rangle$ is far detuned from the other states (e.g., for $T=200$ ns pulses, $|\rm{f}\rangle$ is only populated if $\delta\approx E_{C}/2$ [16]). Thus, after substituting $t_{0}$ in Eq. 10 such that it explicitly depends on $\delta$, we solve for $\delta$ to find the two roots of the threshold condition $p_{\rm{e}}^{\rm{thresh}}=0.99$, denoted $\delta_{-}$ and $\delta_{+}$. Due to symmetry, $\delta_{-}=-\delta_{+}$, and we approximate the detuning robustness as $\Delta_{\rm{rob}}^{\rm{LZ}}=2\delta_{+}$. Combining $\Delta_{\rm{rob}}^{\rm{LZ}}$ with the previously found values of $\max_{t}(p_{\rm{f}})$, we obtain the bounds in Fig. 2: the cyan-colored region for the 50 ns pulses and the magenta-colored region for the 200 ns pulses. These bounds are less stringent than our definition of detuning robustness, as they do not include the constraint that $p_{\rm{e}}(T/2)\geq 0.999$ anywhere between $\delta_{-}$ and $\delta_{+}$. The corresponding limit that does incorporate this additional constraint is shown as the orange dotted boundary, plotted only for the 50 ns pulses, since for the $200$ ns pulses, the limits with and without this constraint nearly coincide. Notably, the region where the orange dotted boundary does not overlap with the cyan-colored boundary indicates that, under the Landau–Zener approximation, the $50$ ns pulses satisfy our stricter definition of detuning robustness only at higher values of $\max_{t}(p_{\rm{f}})$. We note that all models except for those with linear detunings (type-1) and the SECH2 evaluated for $T=50$ ns, which prove less optimal for part of their fronts, lie firmly within these bounds.
@@
Moreover, Fig. 2 reveals that all pulse models are capable of achieving moderate detuning robustness—for instance, $\Delta_{\rm{rob}}=0.1$ GHz—even when the maximum transient population of the second excited state remains low (i.e., $\max_{t}(p_{\rm{f}})$ $\leq 0.01$). Models featuring nonlinear detuning profiles (types 2–4) become essential only when targeting higher levels of detuning robustness. Shorter pulses ($T=50$ ns) can attain modest robustness but typically incur higher maximum transient second-excited state populations. By contrast, comparable—or in some cases greater—detuning robustness can be achieved with reduced $\max_{t}(p_{\rm{f}})$, albeit at the cost of longer pulse durations ($T=200$ ns).
@@
While the Pareto fronts serve as our primary figure of merit, they tell us nothing explicitly about the optimality of the parameters necessary to create them. In fact, each envelope function has a total of three parameters, i.e., $[\Omega_{0},n,\beta\;(\epsilon)]$ for SG (SECH) pulses, and depending on the detuning function, there are anywhere from one (for type-1) to five (type-2) detuning parameters. In the next section, we ascertain the overall importance of each of these parameters with respect to $\max_{t}(p_{\rm{f}})$, detuning robustness $\Delta_{\rm{rob}}$, and amplitude robustness $\Omega_{\rm{rob}}$.
IV Parameter Analysis

### v4

@@ 2.2 Pulse models
where we subtract by $k_{2}$ in Eq. (7) to ensure that we satisfy the condition $\Delta(0)-\delta=0$, which is naturally satisfied for the other detuning functions we consider. For these detuning functions, the parameters $k_{1}$ and $k_{2}$ are searched for in the aforementioned ranges $k_{1}\in[-0.3,0.3]$ GHz and $k_{2}\in[-0.1,0.1]$ GHz; whereas, the parameters $\gamma_{1}$ and $\gamma_{2}$ are each searched for in the range $(0,8]$. The parameter $k_{1}$ acts as a scaling factor which symmetrically affects the value at which the detuning saturates, i.e., the detuning asymptotically approaches $\pm k_{1}$ since $\tanh(x)\rightarrow\pm 1$ when $|\absolutevalue{x}$ is large. However, $\gamma_{1}$ controls how quickly the detuning transitions from negative to positive values, i.e., the steepness of the slope. Small values of $\gamma_{1}$ yield a near-linear detuning and large values yield a sharper transition near $t=0$, with extremely large values producing a step-like detuning. Since Eq. (7) is just Eq. (6) plus an additional term, the effect of these parameters is the same in both detuning functions. Moreover, the additional term in Eq. (7) introduces a dip in the detuning function around $t=0$ and increasing $k_{2}$ makes the dip more pronounced. The parameter $\gamma_{2}$ affects the localization of the dip, with a larger $\gamma_{2}$ yielding a dip more concentrated around $t=0$ and a smaller $\gamma_{2}$ making it spread out over a longer time.
As we have previously introduced the shorthand notations SG and SECH for our two envelope types, we further systematically distinguish our models by referring to Eqs. (4), (5), (6), and (7) as detunings of type 1, type 2, type 3, and type 4, respectively. We then create our model names by agglutinating the SG or SECH suffix to a particular detuning model as shown in Table 1. Thus, in this work, we consider a total of eight models: four with super-Gaussian envelopes and four with hyperbolic secant envelopes. These models each satisfy the conditions $\Omega(0)=\Omega_{0}$, $\Omega(\pm T/2)=0$, $\Delta(0)=\delta$ and are evaluated at pulse durations of $T=50$ ns and $T=200$ ns. These two pulse durations span a range of experimentally relevant gate times in superconducting qubit hardware. Shorter durations are motivated by the need for fast gate operations in near-term devices, whereas longer durations allow greater detuning robustness and reduced leakage into higher excited states. The Pareto-optimal smooth-envelope pulses considered in this work exhibit spectral bandwidths dominated by frequency modulation, instead of the pulse duration itself. Their half-power bandwidth is typically on the order of $\pm 100$ MHz, well within the capabilities of contemporary RF signal generators. As an example, in transmon-based systems, we have previously shown that sufficiently long pulses driven close to resonance with modest modulation depths can exhibit detuning robustness with a negligible final second excited state population, thereby isolating the qubit subspace despite the transmon’s weak anharmonicity [Kuzmanovic_2024].
@@
Model Envelope $\Omega(t)/\Omega_{0}]$ Detuning $\Delta(t)-\delta]$
@@
SG2 $\left(\frac{e^{\beta|2t/T|^{n}}-e^{\beta}}{1-e^{\beta}}\right)$ $\sum\limits_{j=1}^{5}k_{j}(2t/T)^{j}$
@@
Table 1: The two-state-inspired models used in this work, each consisting of a pulse envelope function $\Omega(t)$ and detuning function $\Delta(t)$. Each envelope function has a total of three parameters, i.e., $\Omega_{0},n,\beta\;(\epsilon)]$ for SG (SECH) envelopes. Here, the detuning prefactors are denoted as $k_{1}$, $\cdots$, $k_{5}$, with detuning models of type 1 and 3 only using $k_{1}$, type 4 using $k_{1}$ and $k_{2}$, and type 2 using all 5. The parameters $\gamma_{1}$ and $\gamma_{2}$ only appear in the detuning functions of types 3 and 4. Each model satisfies the conditions $\Omega(0)=\Omega_{0}$, $\Omega(\pm T/2)=0$, and $\Delta(0)=\delta$.
In the following subsection, we discuss the Pareto-optimization technique for finding parameters that, when inserted into the matrix elements of Eq. (1), guide the system to adiabatically follow its lowest eigenstate, thus facilitating robust population transfer from $|\rm{g}\rangle$ to $|\rm{e}\rangle$.
2.3 Pareto-optimization
@@
The conundrum of determining the optimal pulse parameters for our models arises from the fact that a pulse that maximizes detuning robustness $\Delta_{\rm{rob}}$ at a given threshold of $p_{\rm{e}}(T/2)$ will generally not minimize the maximum transient second-excited state population, $\max_{t}(p_{\rm{f}})$, and vice versa. This creates a multiobjective optimization (MOO) problem, where we must balance these competing objectives. To address this trade-off, we apply evolutionary computation methods to identify the best solutions. Techniques such as genetic algorithms and estimation of distribution algorithms have been successfully used in complex optimization problems across various fields, including evacuation planning [Garrett2006] and robotic control optimization [Liang2024]. Genetic algorithms operate by mimicking the principles of natural selection to iteratively evolve a population of candidate solutions toward optimality. The process begins with the generation of an initial population that satisfies predefined bounds and constraints. Each individual in the population is evaluated using an objective function to assign a fitness score. The algorithm then progresses through iterative stages, where parents for the next generation are selected based on their fitness [Miller1995].
@@
In solving this MOO, we rely on the Pareto front, which represents the optimal trade-off between objectives in the objective space. Pareto fronts have been used in multiobservable quantum control, where it was shown that the smoothness and convexity of the Pareto front depend on the commutativity of the observables [Sun2017]. To find the Pareto front numerically, we use MATLAB’s gamultiobj() function, which implements a controlled, elitist genetic algorithm based on the non-dominated sorting genetic algorithm (NSGA-II) [Deb2002].
@@
Building on these general principles, NSGA-II enhances the process with additional layers of sophistication, such as non-dominated sorting and crowding distance measures. After scoring the offspring based on their objective function values and feasibility, the current population and offspring are merged into a single set. This set is sorted into non-dominated fronts and individuals are ranked based on both their Pareto rank and their crowding distance—a measure of solution diversity within a front. Finally, the population is trimmed to its original size, retaining individuals from higher-ranking fronts and preserving diversity. This iterative process continues until convergence criteria are met, yielding an accurate approximation of the Pareto front and enabling efficient exploration of multiobjective optimization spaces.
3 Results
@@
In this section, we present the Pareto fronts for each of our models which optimally minimizes the maximum transient second-state population $\max_{t}(p_{\rm{f}})$ in the range $t\in[-T/2,T/2]$ given that we must simultaneously maximize the detuning robustness at the threshold $p_{\rm{e}}(T/2)\geq 0.99$. These fronts are evaluated for pulse durations of $T=50$ ns and $T=200$ ns. Since this threshold may be reached at multiple points across the detuning (or amplitude) axis, a consistent definition of robustness is essential. We therefore define detuning robustness (or amplitude robustness) as the width of the $p_{\rm{e}}(T/2)$ curve between the first two points where $p_{\rm{e}}(T/2)=0.99$, provided that $p_{\rm{e}}(T/2)\geq 0.999$ is reached at least once within this interval. We include amplitude robustness in parenthesis, because even though it was not one of the objectives in our MOO problem, we compute it using the optimal pulse parameters found by the genetic algorithm.
@@
Moreover, the Pareto fronts for each of our models at the two pulse durations $T$ are presented in Fig. 2(a). Here, we observe that the models form two distinct clusters: one for $T=50$ ns pulses (solid curves) and one for $T=200$ ns pulses (dashed curves), with the latter affording us more detuning robustness. For detuning robustness values up to approximately $0.1$ GHz for 50 ns pulses and $0.15$ GHz for 200 ns pulses, all models yield similar $\max_{t}(p_{\rm{f}})$, with the clusters corresponding to respective pulse durations. This suggests that, within these moderate robustness regimes, performance is largely insensitive to the specific choice of pulse model.
@@
We further note that the models with linear detuning functions perform less optimally compared to the other detuning types at higher detuning robustness levels, as evidenced by the divergence towards higher $\max_{t}(p_{\rm{f}})$ values at approximately 0.13 GHz for $50$ ns pulses and at approximately 0.22 GHz for $200$ ns pulses. Otherwise, it is generally less obvious how the other models fare relative to each other. Thus, to better contextualize these results, we assign approximate theoretical bounds to these fronts to ascertain how close we are to the limits imposed by adiabaticity.
@@
Figure 2: (a) The maximum transient second excited state population $\max_{t}(p_{\rm{f}})$ vs detuning robustness for various models with pulse durations of 50 ns (solid curves) and 200 ns (dashed curves). The shaded regions (cyan and magenta for 50 ns and 200 ns pulses, respectively) represent the range of data constituting the fronts found from using the Landau-Zener (LZ) formula to approximate the theoretical limit of the detuning robustness for each model. These LZ-based bounds are computed at the threshold condition $p_{\rm{e}}^{\rm{thresh}}=0.99$, but without requiring that $p_{\rm{e}}(T/2)\geq 0.999$ anywhere within the corresponding detuning interval (i.e., between the roots $\delta_{-}$ and $\delta_{+}$, where $p_{\rm{e}}(\delta_{\pm})=0.99$). The orange dashed boundary includes this additional constraint. As an example, panel (b) shows the time profiles of the envelope function and detuning function for the SECH3 model at maximum detuning robustness. The upper (lower) tile corresponds to $T=50$ ($200$) ns.
@@
To this end, we use the Landau-Zener (LZ) formula in calculating the detuning robustness at the threshold $p_{\rm{e}}(T/2)\geq 0.99$. The LZ formula provides the diabatic transition probability:
@@
where $\Gamma=\Omega(t_{0})^{2}/4\hbar|v|$ with $t_{0}$ defined as the time at which the detuning vanishes, $\Delta(t_{0})=0$. In the limit $\Gamma\gg 1$, the transition is adiabatic, whereas for $\Gamma\ll 1$, it is nonadiabatic. Here, $v$ is the Landau-Zener velocity, representing the slope of $\Delta(t)$ at the avoided crossing:
$v=\left.\frac{d\Delta(t)}{dt}\right|_{t=t_{0}}.$ (9)
For example, in the case of a linear detuning $\Delta(t)=vt+\delta$, we have $t_{0}=-\delta/v$.
In terms of $v$ and considering the threshold value of the first excited state population we use for calculating the detuning robustness, i.e., $p_{\rm{e}}^{\rm{thresh}}=0.99$, we impose the condition
@@
Here, we subtract from unity as Eq. (8) describes the probability of remaining in $|\rm{g}\rangle$ at the level crossing and we further assume that $|\rm{f}\rangle$ is far detuned from the other states (e.g., for $T=200$ ns pulses, $|\rm{f}\rangle$ is only populated if $\delta\approx E_{C}/2$ [Kuzmanovic_2024]). Thus, after substituting $t_{0}$ in Eq. (10) such that it explicitly depends on $\delta$, we solve for $\delta$ to find the two roots of the threshold condition $p_{\rm{e}}^{\rm{thresh}}=0.99$, denoted $\delta_{-}$ and $\delta_{+}$. Therefore, $\Delta_{\rm{rob}}^{\rm{LZ}}=\delta_{+}-\delta_{-}$.
These bounds are less stringent than the requirement on the transfer fidelity imposed on the Pareto-optimal solutions, as they do not include the constraint that $p_{\rm{e}}(T/2)\geq 0.999$ anywhere between $\delta_{-}$ and $\delta_{+}$. The corresponding limit that does incorporate this additional constraint is shown as the orange dashed boundary, discernible only for the 50 ns pulses, since for the $200$ ns pulses, the limits with and without this constraint nearly coincide.
[…]
Notably, the $50$ ns pulses do not fall within the bounds obtained based on an adiabatic Landau-Zener trajectory, save for the ones with low detuning robustness. In contrast, the opposite is true for the $200$ ns pulses, where they operate within the adiabatic regime with the exception of the linearly-detuned models at extreme detuning robustness.
@@
Nonetheless, Fig. 2(a) reveals that all pulse models are capable of achieving moderate detuning robustness—for instance, $\Delta_{\rm{rob}}=0.1$ GHz—even when the maximum transient population of the second excited state remains low (i.e., $\max_{t}(p_{\rm{f}})$ $\leq 0.01$). Models featuring nonlinear detuning profiles (types 2–4) become essential only when targeting higher levels of detuning robustness. Shorter pulses ($T=50$ ns) can attain modest robustness but typically incur higher maximum transient second-excited state populations. By contrast, comparable—or in some cases greater—detuning robustness can be achieved with reduced $\max_{t}(p_{\rm{f}})$, albeit at the cost of longer pulse durations ($T=200$ ns).
@@
While the Pareto fronts serve as our primary figure of merit, they do not explicitly reveal the values of the Pareto-optimal parameters. In fact, each envelope function has a total of three parameters, i.e., $\Omega_{0},n,\beta\;(\epsilon)]$ for SG (SECH) pulses and depending on the detuning function, there are anywhere from one (for type-1) to five (type-2) detuning parameters. As an example, Fig. 2(b) shows the envelope and detuning functions at $T=50$ ns (upper tile) and at $T=200$ ns (lower tile) for the SECH3 model when its detuning robustness is greatest in our Pareto solution. The parameters which generate these functions for $T=50$ ($200$) ns are $\Omega_{0}=0.1$ ($0.0582$) GHz, $n=0.995$ ($1.14$), $\epsilon=2$ ($3.15$), $k_{1}=-0.161$ (0.152)(-0.152) GHz, and $\gamma_{1}=1.92$ $(2.92)$. In the next section, we ascertain the overall importance of each model’s parameters with respect to $\max_{t}(p_{\rm{f}})$, detuning robustness $\Delta_{\rm{rob}}$, and amplitude robustness $\Omega_{\rm{rob}}$.
4 Parameter Analysis